Thien has quarters and nickles in his pocket. he has ten fewer nickles than quarters, and their total value is $6.10. how many of each does he have?
Answer by mananth(16946)

quarters 25 x
nickels 5 (x-10)
Value= 610
25x+5x-50= 610
30x -50=610
30x=610
30x=660
/30
x=22 quarters 550
12 Nickels 60